# Real-Time Simulated Avatar from Head-Mounted Sensors

### v1

@@ 3.3 Simulated Humanoid Control From Head-Mounted Sensors
For the 6DoF headset pose ${\bm{{q}}_{t}}^{\tau}$, we treat it as a virtual “joint”, where $\bm{q}^{\tau}_{t+1}\triangleq(\bm{p}^{\tau}_{t+1},\bm{\theta}^{\tau}_{t+1})$ contains the global rotation $\bm{\theta}^{\tau}_{t+1}$ and translation $\bm{p}^{\tau}_{t+1}$ of the headset. We use the humanoid head to track the pose of the headset by computing the rotation and translation difference between the head joint ${\bm{{q}}_{t}}^{\text{Head}}$ and the headset pose $\bm{q}^{\tau}_{t+1}$. This creates the headset pose feature: $\bm{\psi}_{t}=({\bm{{\theta}}_{t}}^{\text{Head}}\ominus\bm{\theta}^{\tau}_{t+1},{\bm{{p}}_{t}}^{\text{Head}}-\bm{p}^{\tau}_{t+1})$. The image feature and the headset pose features are then concatenated with proprioception ${\bm{s}^{\text{p}}_{t}}$ to form the input to an MLP to compute joint actions, as illustrated in Fig.3.
@@ 4 Experiments
Table 2: Pose estimation result on the test split ($458$k frames) of synthetic data and real-world captures ($40$k frames). Here, our MPJPE is computed as “device-relative” instead of root-relative.
@@
[…]
@@
We adopt the SOTA vision-based pose estimation method UnrealEgo [5] as the main vision-based baseline. UnrealEgo uses a U-Net structure to first reconstruct 2D heatmaps from input images. Then, an autoencoder is used to lift the 2D heatmap to 3D keypoints. UnrealEgo predicts the 3D pose in the headset’s coordinate system instead of the global one. To compare against a physics-based method, we reimplement KinPoly [30] and add image input to create KinPoly-v. We also remove its dependence on action labels and external forces (replacing UHC [30] with PHC [33]). KinPoly-v uses a two-stage method: first estimate full-body pose from images, and then feed them into a pretrained imitator to drive the simulated avatar. It uses a closed-loop system, where the simulated state is fed into the image-based pose estimator, making it “dynamically regulated”. KinPoly-v shares the same overall architecture as our method but uses kinematic pose as an intermediate representation to communicate with a pretrained imitator, instead of an end-to-end approach.
@@ C.1 Synthetic Data Generation
Table A.1: Error analysis based on joint in-frame status. “In-frame” is not equivalent to “visible” due to self-occlusion.
@@
Here we conduct a visibility analysis on our generated synthetic data. As accounting for self-occlusion involves reprocessing the synthetic dataset and conducting ray-marching for each joint on the clothed mesh to ascertain visibility, we opt to use the “in-frame” (whether the joint is in one of the camera frames) statistics to approximate visibility. This measure is a reliable visibility indicator for upper body joints, but less so for the lower body, where self-occlusion and extreme viewpoints are more prevalent (see Figure A.2). From Table A.1 we can see that the shoulder and elbow joints are frequently outside the frame, while the wrist and leg joints often remain within the frame. For the shoulder, elbow, and knee joints, their in-frame and out-frame results are similar, since they are closer to the torso. For the body extremities (wrists and ankles), there is a clear gap, as the largest errors occur out of the frame.

### v2

@@ 3.3 Simulated Humanoid Control From Head-Mounted Sensors
For the 6DoF headset pose ${\bm{{q}}_{t}}^{\tau}$, we treat it as a virtual “joint", where $\bm{q}^{\tau}_{t+1}\triangleq(\bm{p}^{\tau}_{t+1},\bm{\theta}^{\tau}_{t+1})$ contains the global rotation $\bm{\theta}^{\tau}_{t+1}$ and translation $\bm{p}^{\tau}_{t+1}$ of the headset. We use the humanoid head to track the pose of the headset by computing the rotation and translation difference between the head joint ${\bm{{q}}_{t}}^{\text{Head}}$ and the headset pose $\bm{q}^{\tau}_{t+1}$. This creates the headset pose feature: $\bm{\psi}_{t}=({\bm{{\theta}}_{t}}^{\text{Head}}\ominus\bm{\theta}^{\tau}_{t+1},{\bm{{p}}_{t}}^{\text{Head}}-\bm{p}^{\tau}_{t+1})$. The image feature and the headset pose features are then concatenated with proprioception ${\bm{s}^{\text{p}}_{t}}$ to form the input to an MLP to compute joint actions, as illustrated in Fig.3.
@@ 4 Experiments
Table 2: Pose estimation result on the test split ($458$k frames) of synthetic data and real-world captures ($40$k frames). Here, our MPJPE is computed as “device-relative" instead of root-relative.
@@
We report both pose- and physics-based metrics to evaluate our avatar’s performance. We report the success rate (Succ) as in UHC [30], defined as: at every point during imitation, the head joint is $<0.5$m from the headset pose. For pose estimation, we report the device-relative (instead of root relative) per-joint position error (MPJPE) $E_{\text{mpjpe}}$, global MPJPE $E_{\text{g-mpjpe}}$, and MPJPE after Procrustes analysis $E_{\text{pa-mpjpe}}$. Since $E_{\text{pa-mpjpe}}$ solves for the best matching scale, rotation, and translation, it is more suitable for our real-world dataset, where the scale and global position of the pseudo-ground-truth pose are noisy. To maintain the consistency of evaluation between the two humanoids, we select 11 common joints (head, left and right shoulders, elbows, wrists, knees, ankles) to report joint errors, rather than evaluating all joints as in the prior art [33]. To test physical realism, we include acceleration $E_{\text{acc}}$ (mm/frame2) and velocity $E_{\text{vel}}$ (mm / frame) errors.
@@
We adopt the SOTA vision-based pose estimation method UnrealEgo [5] as the main vision-based baseline. UnrealEgo uses a U-Net structure to first reconstruct 2D heatmaps from input images. Then, an autoencoder is used to lift the 2D heatmap to 3D keypoints. UnrealEgo predicts the 3D pose in the headset’s coordinate system instead of the global one. To compare against a physics-based method, we reimplement KinPoly [30] and add image input to create KinPoly-v. We also remove its dependence on action labels and external forces (replacing UHC [30] with PHC [33]). KinPoly-v uses a two-stage method: first estimate full-body pose from images, and then feed them into a pretrained imitator to drive the simulated avatar. It uses a closed-loop system, where the simulated state is fed into the image-based pose estimator, making it “dynamically regulated". KinPoly-v shares the same overall architecture as our method but uses kinematic pose as an intermediate representation to communicate with a pretrained imitator, instead of an end-to-end approach.
@@ Conclusions
Acknowledgement
We thank Zihui Lin for her help in making the plots in this paper. Zhengyi Luo is supported by the Meta AI Mentorship (AIM) program.
@@ C.1 Synthetic Data Generation
Table A.1: Error analysis based on joint in-frame status. “In-frame" is not equivalent to “visible" due to self-occlusion.
@@
Here we conduct a visibility analysis on our generated synthetic data. As accounting for self-occlusion involves reprocessing the synthetic dataset and conducting ray-marching for each joint on the clothed mesh to ascertain visibility, we opt to use the “in-frame" (whether the joint is in one of the camera frames) statistics to approximate visibility. This measure is a reliable visibility indicator for upper body joints, but less so for the lower body, where self-occlusion and extreme viewpoints are more prevalent (see Figure A.2). From Table A.1 we can see that the shoulder and elbow joints are frequently outside the frame, while the wrist and leg joints often remain within the frame. For the shoulder, elbow, and knee joints, their in-frame and out-frame results are similar, since they are closer to the torso. For the body extremities (wrists and ankles), there is a clear gap, as the largest errors occur out of the frame.